JS has decided to purchase t-shirts and print them with a logo to commemorate a major international sporting event.

Sales

The commemorative t-shirts will be sold for $10 each and predicted sales are as follows:

July                  9,000 t-shirts

August             18,000 t-shirts

September        22,500 t-shirts

One third of sales will be for cash. The remainder will be on credit with the customer paying the month after sale.

Purchases

The t-shirts will cost $6 each and will be purchased in the month prior to sale. It is expected that 10% of the t-shirts purchased will be damaged during the printing process and will not be suitable for sale. The supplier has offered two months credit.

Capital investment

To print the t-shirts with the logo of the sporting event will require the purchase of a machine costing $30,000. The machine will be bought at the start of the project and paid for in August. The machine will have a five-year useful life but no expected residual value. The machine will be used elsewhere in the business at the end of this project.

Expenses

Expenses, excluding advertising, of $20,000 per month will be incurred each month and paid in the month incurred. Advertising costs of $5,000 per month will be incurred in each of the months July, August and September and will be paid one month in arrears.

Required:

Produce a cash budget for the project for each of the three months July, August and September.
Answer #1

July

Cash income = cash sales of July = (9000 t-shits*90%*$10*(1/3))= $27000

Cash expenses

Cash expenses = $ 20,000

Net cash income = $(27000-20000)= $7000

Aug

Cash income

a) Cash sales of Aug = (18000*90%*(1/3)*$10) = $ 54000

b) Cash receivable for July credit sales = (9000*90%*(2/3)*$10) = $ 54000

Total Cash income = $108000

Cash Expense

a) Purchase price of 9000 t-shirts in July = (9000*$6) = $ 54000

b) Machine purchase payment = $ 30000

c) Advertisement cost of July = $ 5000

d) Other expenses = $20000

Total Cash expenses = a+b+c+d = $ 109000

Thus net cash loss of $ 1000.

Sept

Cash Income

a) Cash sales for Sept = (22500*90%*(1/3)*$10) = $67500

b) Cash received for sales in Aug = (18000*90%*(2/3)*$10) = $ 108000

Total Cash income = a+b = $175500

Cash expenses

a) Other expenses = $ 20000

b) advertisement cost of Aug = $ 5000

c) Purchase price of 18000 t-shirts of Aug = (18000*$6) = $108000

Total cash expenses = a+b+c = $ 133000

Thus net cash income = $ 42500